how do you find the discriminant of each quadratic equation then state the number and type of solutions of 14x^2-10x+21=14?
Answer by stanbon(75887)

---
find the discriminant of each quadratic equation then state the number and type of solutions of 14x^2-10x+21=14?
Rearrange:
14x^2 - 10x + 7 = 0
-----
a = 14 ; b = -10, c = 7
---
Discriminant: b^2 - 4ac = (-10)^2 -4*14*7 = -292
--------------------
Since the discriminant is negative, there are two Complex Number solutions.
---------
Note: If the discriminant had been zero, there would be one Real Number solution
of multiplicity 2.
If the discriminant had been positive, there would be two Real Number solutions.
==============================
Cheers,
Stan H.